OpenStudy (anonymous):

A rock have 3N of weight force, tied to a string of 2.5 m long, describes a horizontal circle of 2 m radius. The twine, fixed at one end, generates a conical surface. Determine: a) The tensile strength of the wire, in newtons; b) the frequency of rotation in Hz consider g = 10 m / s ^ 2

OpenStudy (anonymous):

here tension in the string provides req Centripetaal force

OpenStudy (anonymous):

I coud to find the tension it's 5N, but the frequency of rotation is very hard

OpenStudy (anonymous):

v = 1/T

OpenStudy (anonymous):

you can find f by solve T

OpenStudy (anonymous):

why v= 1/T ?

OpenStudy (anonymous):

can you get velocity the rock goes around

OpenStudy (anonymous):

how? with that formula?

OpenStudy (anonymous):

you did find T right? then sum of rock's weight and tension must be central force.

OpenStudy (anonymous):

then you can get velocity by F=(mv^2)/r

OpenStudy (anonymous):

I din't know this formula

OpenStudy (anonymous):

why?? it's the force that is acting on object when the object moves circular motion

OpenStudy (anonymous):

so it's more easy that I thought, the ask is = 0.4 thank you

OpenStudy (anonymous):